OpenStudy (anonymous):

Evaluate: x^-6/y^-5 * y^-5/x^-6

OpenStudy (anonymous):

please rewrite it clearer ... I don't get it at all ! O_O :(

OpenStudy (anonymous):

is it? ...\[\Large \frac{x^{-6}}{y^{-5}}\cdot \frac{y^{-5}}{x^{-6}}\]

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

ok.. sweet, then this is what we're going to do... Use this formula... \[\LARGE \frac{a^m}{a^n}=a^{m-n}\]

OpenStudy (anonymous):

i can't use that formula because there are two different denominators

OpenStudy (anonymous):

i mean two different bases

OpenStudy (anonymous):

we have...\[\LARGE \frac{x^{-6}\cdot y^{-5}}{x^{-6}\cdot y^{-5}}=x^{-6-(-6)}\cdot y^{-5-(-5)}\]

OpenStudy (anonymous):

from there you get... ??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Nice ;)